a rectangular lot measures 66m by 107m the sidewalk is 2m wide what is the area of the sidewalk
Answer by checkley77(12844)

66*107=7,062 m^2 is the lot area.
(66+4)(107+4)
70*111=7,770 is the total area including the sidewalk.
7,770-7,062=708m^2 is the area of the sidewalk.